Lilian has part of $9000 invested at 4% and the remainder at 6%. His annual return on each investment is the same. At what rate would she have to invest all her money to get the same interest?
PLEASE USE SYSTEMS OF EQUATIONS AND SHOW YOUR WORK !!!

Answer by addingup(3677)

Let x be the amount invested at 4%
0.04x+0.06(9,000-x) = annual return
0.04x+54-0.06x = annual return
-0.02x = -54
x = 2,700 This is the amount invested at 4%
9,000-2,700 = 6,300 is the amount invested at 6%
````````````````````````````````
2700/9000 = 30% is invested at 4%
100-30 = 70% is invested at 6%
0.30*0.04 = 0.012
+
0.70*0.06 = 0.042
------------------
Blended interest: 0.054
To get the same interest she'd have to invest all her money at 5.4%